OpenStudy (anonymous):

Find the zeros of the following quadratic functions, which are already in factored form: F(x)= (700-5x)(50+2x) Can someone help walk me through the steps ?

OpenStudy (campbell_st):

to find the zeros set each binomial to zero and solve for x so you need to solve 700 - 5x = 0 and 50 + 2x = 0 the reason there are called zeros, is that then either of the solutions above are substituted into the original equation yo'll get F(x) = 0

OpenStudy (anonymous):

Okay.. but I don't get how to find the zeros.

OpenStudy (campbell_st):

well you just solve the equations for 700 - 5x = 0 start by subtracting 5x from both sides of the equation 700 = 5x now divide both sides by 5 140 = x so x = 140 is a zero do the same for the other binomial.. 50 +2x = 0 what value of x makes the equation equal zero... hope this helps

OpenStudy (anonymous):

So it's 25 = x for the next one ?

OpenStudy (campbell_st):

nope its x = -25